# A fractal geometry immersed in a hierarchical magnetic flux distribution

### v1

@@ I Introduction
Inspired by these experimental advancements in creating the fractal structures in reality, various interesting theoretical studies concerning the deterministic fractal structures have taken place in recent times. This includes works on energy level statistics in fractal clusters yorikawa-JPC2019 ; askar-PRB2019 , confinement of electrons in fractal plane using electric field yuan-PRB2020 , electronic fractal pattern in building molecular fractal systems latge-pccp2022 , role of anisotropy on magneto-transport in fractal network maiti-PLA2020 , and Josephson effect in a fractal geometry vladimir-arXiv2024 among others.
@@
It is expected that, in future, further advancement in nanotechnology and fabrication techniques will enable us to do more delicate experiments with these fractal lattice models. Such possibility inspires us to undertake this theoretical investigation presented in this article. We consider a Sierpinski gasket (SPG) fractal lattice structure immersed in a deterministic hierarchical distribution of magnetic fluxes. The ramification of a SPG fractal structure is such that, it offers us with down-headed triangular loops of gradually increasing size (See Fig. 1). Following this ramification pattern, we introduce a hierarchical distribution of magnetic fluxes $\Phi_{\ell}=\lambda^{\ell}\Phi$ ($\lambda$ being the hierarchy parameter) through these down-headed triangular closed loops. Such distribution of the magnetic fluxes will be manifested as a distribution of Aharonov-Bohm (AB) phases, whenever an electron will hop from one atomic site to another atomic site along these down-headed triangular closed loops. This provides us with an interesting situation, where we have a deterministic disorder in the distribution of the AB phases in this closed-loop fractal geometry controlled by a hierarchy parameter $\lambda$.
@@ III Effect of the hierarchy parameter on the energy eigenvalue spectrum
In this section, we discuss the role of the hierarchy parameter $\lambda$ on the energy eigenvalue spectrum of the SPG lattice. We have first explicitly constructed the Hamiltonian matrix for the SPG lattice using Eq. (2). We take the on-site potential $\epsilon_{n}=0$ for all the sites, and consider only nearest-neighbor constant coupling between the sites $t_{mn}=t=1$. The direction of the hierarchical fluxes $\Phi_{\ell}$ (where $\ell=1,2,3,\ldots$) in different down-headed triangular closed plaquettes is taken to be in the out-of-plane direction. Accordingly, the electrons will pick up the AB phases ±$\pm\theta_{\ell}$ (‘$+$’ is for forward hopping and ‘--’ is for backward hopping). Once the Hamiltonian matrices are constructed for different generations of the SPG lattice, one can easily diagonalize them to evaluate the energy eigenvalue spectrum for the lattice structure. Here, we present the results for the 3rd and the 4th generation SPG lattice structure in Fig. 2 and Fig. 3, respectively.
Figure 2: Variation of the energy eigenvalue spectrum for a 3rd generation SPG lattice structure [with $\mathcal{N}(3)=42$] as a function of the hierarchical magnetic flux for different values of the hierarchy parameter, viz., (a) $\lambda=0$, (b) $\lambda=0.1$, (c) $\lambda=0.5$, and (d) $\lambda=0.9$.
@@
Figure 3: Energy eigenvalue distribution for a 4th generation SPG lattice structure [with $\mathcal{N}(4)=123$] as a function of the hierarchical magnetic flux with values of the hierarchy parameter as (a) $\lambda=0$, (b) $\lambda=0.1$, (c) $\lambda=0.5$, and (d) $\lambda=0.9$.
@@
Figure 5: Variation of the persistent current ($I_{\textrm{max}}$) as a function of the hierarchy parameter $\lambda$ for a 3rd generation SPG fractal lattice. (a) is for a filling factor $\nu=1/3$, (b) is for $\nu=1/2$, and (c) represents $\nu=2/3$.
@@
It is important to see the manifestation of these properties in terms of some experimentally measurable quantities. This brings us to the next section, where we demonstrate how one can control the persistent current as a function of the hierarchy parameter in this closed-loop fractal geometry.
@@ IV Controlling the persistent current
where $c$ is a constant which is set equal to 1 for our calculation. The total persistent current in the system is given by,
@@
For our model, the magnetic flux $\Phi$ depends on the hierarchy parameter $\lambda$. So, we can study the magnetic response of the SPG fractal geometry in terms of the persistent current as a function of $\lambda$. We have systematically investigated the behavior of the persistent current in a 3rd generation SPG fractal network as a function of the hierarchy parameter $\lambda$ for different values of $\nu$. This is illustrated in Fig. 5(a)-(c). It is observed that, for different filling factors $\nu$, the persistent current in the system gets enhanced as one increases the value of $\lambda$. This opens up an interesting possibility by which one can flexibly control the persistent current in such multi-loop fractal geometries. Once again, this feature is also tested for a 4th generation SPG fractal structure (Fig. 6), and found to be true as an evidence of the generic self-similar property of a fractal structure.
@@ V Summary and future outlook
In conclusion, we have studied the role of a hierarchical distribution of magnetic flux on the energy spectrum and persistent current in a multi-loop Sierpinski gasket fractal structure. We have shown that, by controlling the hierarchy parameter of the embedded magnetic flux in this SPG fractal network, one can effectively engineer quantum states in such a complex system. From the application point-of-view, we have shown that this directly impacts the behavior of the persistent current in this SPG fractal structure. We can suitably control the persistent current in the system by tunning the hierarchy parameter in a flexible way for different filling factors of the electronic states. One could in principle, test these results by designing an experiment using a self-assembled molecular SPG fractal structure. This could be helpful to devise an unconventional nanoelectronic device using a molecular SPG fractal structure. In future, we plan to generalize this idea for other similar kind of closed-loop fractal geometries, such as Vicsek fractal pal-PRB2012 ; pal-PRA2013 .
@@ Acknowledgements.
The author would like to thank Prof. Arunava Chakrabarti from Presidency University, Kolkata for his useful feedback on the initial stage of this work. The author also acknowledges helpful discussion with Amrita Mukherjee from TIFR, Mumbai on the calculation of the persistent current.

### v2

@@ I Introduction
Inspired by these experimental advancements in creating the fractal structures in reality, various interesting theoretical studies concerning the deterministic fractal structures have taken place in recent times. This includes works on energy level statistics in fractal clusters yorikawa-JPC2019 ; askar-PRB2019 , confinement of electrons in fractal plane using electric field yuan-PRB2020 , electronic fractal pattern in building molecular fractal systems latge-pccp2022 , role of anisotropy on magneto-transport in fractal network maiti-PLA2020 , Josephson effect in a fractal geometry vladimir-APL2024 , and stability of loop current states in fractal structures built with Bose-Einstein condensates anna-PRA2024 among others.
@@
It is expected that, in future, further advancement in nanotechnology and fabrication techniques will enable us to do more delicate experiments with these fractal lattice models. Such possibility inspires us to undertake this theoretical investigation presented in this article.
We consider a Sierpinski gasket (SPG) fractal lattice structure immersed in a deterministic hierarchical distribution of magnetic fluxes. The ramification of a SPG fractal structure is such that, it offers us with down-headed triangular loops of gradually increasing size (See Fig. 1). Following this ramification pattern, we introduce a hierarchical distribution of magnetic fluxes $\Phi_{\ell}=\lambda^{\ell}\Phi$ ($\lambda$ being the hierarchy parameter) through these down-headed triangular closed loops. Such distribution of the magnetic fluxes will be manifested as a distribution of Aharonov-Bohm (AB) phases, whenever an electron will hop from one atomic site to another atomic site along these down-headed triangular closed loops. This provides us with an interesting situation, where we have a deterministic disorder in the distribution of the AB phases in this closed-loop fractal geometry controlled by a hierarchy parameter $\lambda$. It is to be noted that, in Fig. 1, although schematically we have shown that the fluxes are embedded in the downward triangles, but in reality, along the arms of the upward triangles, we will also have various distribution of the Aharonov-Bohm phases depending on their position in the fractal geometry for any given finite generation. This means that, the upward triangles are also containing some effective fluxes. One may realize such a complex distribution of magnetic fluxes by putting up some synthetic gauge fields gardin-arXiv2024 in various downward plaquettes, which may be controlled locally in each plaquette.
@@ III Effect of the hierarchy parameter on the energy eigenvalue spectrum
In this section, we discuss the role of the hierarchy parameter $\lambda$ on the energy eigenvalue spectrum of the SPG lattice. We have first explicitly constructed the Hamiltonian matrix for the SPG lattice using Eq. (2). We take the on-site potential $\epsilon_{n}=0$ for all the sites, and consider only nearest-neighbor constant coupling between the sites $t_{mn}=t=1$. The direction of the hierarchical fluxes $\Phi_{\ell}$ (where $\ell=1,2,3,\ldots$) in different down-headed triangular closed plaquettes is taken to be in the out-of-plane direction. Accordingly, the electrons will pick up the AB phases ±$\pm\theta_{\ell}$ (‘$+$’ is for forward hopping and ‘--’ is for backward hopping). Once the Hamiltonian matrices are constructed for different generations of the SPG lattice, one can easily diagonalize them to evaluate the energy eigenvalue spectrum for the lattice structure. Here, we present the results for the 3rd and the 4th generation SPG lattice structure in Fig. 2 and Fig. 3, respectively. We note that, the energy ($E$) of the particle is measured in units of the hopping integral $t$.
Figure 2: Variation of the energy eigenvalue spectrum for a 3rd generation SPG lattice structure (with $\mathcal{N}(3)=42$) as a function of the hierarchical magnetic flux for different values of the hierarchy parameter, viz., (a) $\lambda=0$, (b) $\lambda=0.1$, (c) $\lambda=0.5$, and (d) $\lambda=0.9$.
@@
Figure 3: Energy eigenvalue distribution for a 4th generation SPG lattice structure (with $\mathcal{N}(4)=123$) as a function of the hierarchical magnetic flux with values of the hierarchy parameter as (a) $\lambda=0$, (b) $\lambda=0.1$, (c) $\lambda=0.5$, and (d) $\lambda=0.9$.
@@
Figure 5: Variation of the persistent current ($I_{\textrm{max}}$) as a function of the hierarchy parameter $\lambda$ for a 3rd generation SPG fractal lattice for the temperature $T=0$ and $T>0$, respectively. (a) is for a filling factor $\nu=1/3$, (b) is for $\nu=1/2$, and (c) represents $\nu=2/3$.
@@
It is important to see the manifestation of the above properties in terms of some experimentally measurable quantities. This brings us to the next section, where we demonstrate how one can control the persistent current as a function of the hierarchy parameter in this closed-loop fractal geometry.
@@ IV Controlling the persistent current
where $c$ is a constant which is set equal to 1 for our calculation. We note that, for our model, we have the hierarchical distribution of the magnetic fluxes as $\Phi_{\ell}=\lambda^{\ell}\Phi$, where $\lambda$ is the hierarchy parameter, $\ell$ is the generation index, and $\Phi$ is some fundamental flux which can take values from $-\Phi_{0}$ to $\Phi_{0}$. Therefore, $\Phi$ in Eq. (4), for our model can be expressed in terms of various orders of $\Phi_{\ell}$ and $\lambda^{\ell}$ depending on any given ‘$\ell$’-th generation fractal system. The total persistent current in the system is given by,
@@
For our model, the magnetic flux $\Phi$ depends on the hierarchy parameter $\lambda$. So, we can study the magnetic response of the SPG fractal geometry in terms of the persistent current as a function of $\lambda$. We have systematically investigated the behavior of the persistent current in a 3rd generation SPG fractal network as a function of the hierarchy parameter $\lambda$ for different values of $\nu$. This is illustrated in Fig. 5(a)-(c). It is observed that, for different filling factors $\nu$, the persistent current in the system gets enhanced as one increases the value of $\lambda$. This opens up an interesting possibility by which one can flexibly control the persistent current in such multi-loop fractal geometries. Once again, this feature is also tested for a 4th generation SPG fractal structure (Fig. 6), and found to be true as an evidence of the generic self-similar property of a fractal structure. We remark that, in a real experimental setup, if one typically chooses the system size to be $\sim\mu$m, magnetic field $\sim 10$ Tesla, then the persistent current we get will be $\sim$ nano-Amp (nA).
@@
It is important to briefly mention the effect of the finite temperature on the persistent current for our fractal lattice system. At finite temperature ($T>0$), one can calculate the persistent current in the system using the following formula bloch-PRB1970 ; wendler-PSSB1995 :
$I(\Phi)=-\dfrac{\partial F(\Phi)}{\partial\Phi}$ (6a)
with $F(\Phi)=-k_{B}T\>\ln\Big{(}\sum_{\nu}-\dfrac{E_{\nu}}{k_{B}T}\Big{)}$ (6b)
where $F$ is the Helmholtz free energy, $k_{B}$ is the Boltzmann constant, and $T$ is the absolute temperature. It is found that, at finite temperature, the persistent current in the system decreases drastically as evident from Fig. 5 and Fig. 6. This is because, as we go to the finite temperature, electrons can occupy the higher energy levels with different phases, which tend to cancel out the net current in the system.
Figure 7: Variation of the persistent current ($I$) as a function of the magnetic flux for three different values of the hierarchy parameter $\lambda=0.3$ (red solid line), $\lambda=0.6$ (blue dashed line), and $\lambda=0.9$ (green dotted line), respectively in a $3$rd generation SPG fractal structure. The filling factor is set be $\nu=1/3$.
The variation of the persistent current with the magnetic flux for different fixed values of the hierarchy parameter $\lambda$ is shown in Fig. 7. From these plots, we can clearly identify that as the value of $\lambda$ increases, the magnitude of the persistent current in the system also increases. This result is consistent with the result shown in Fig. 5. We note that, the periodicity of the persistent current as a function of the magnetic flux also increases as we increase the value of $\lambda$. This is because of the hierarchical nature of the magnetic flux distribution in the system as a function of the hierarchy parameter $\lambda$.
@@ V Summary and future outlook
In conclusion, we have studied the role of a hierarchical distribution of magnetic flux on the energy spectrum and persistent current in a multi-loop Sierpinski gasket fractal structure. We have shown that, by controlling the hierarchy parameter of the embedded magnetic flux in this SPG fractal network, one can effectively engineer quantum states in such a complex system. From the application point-of-view, we have shown that this directly impacts the behavior of the persistent current in this SPG fractal structure. We can suitably control the persistent current in the system by tunning the hierarchy parameter in a flexible way for different filling factors of the electronic states. One could in principle, test these results by designing an experiment using a self-assembled molecular SPG fractal structure. This could be helpful to devise an unconventional nanoelectronic device using a molecular SPG fractal structure. In future, we plan to generalize this idea for other similar kind of closed-loop fractal geometries, such as Vicsek fractal pal-PRB2012 ; pal-PRA2013 . Another interesting future prospect of this work could be to study these properties in fractal lattices with self-affine feature self-affine .
@@ Acknowledgements.
The author would like to thank Prof. Arunava Chakrabarti from Presidency University, Kolkata for his useful feedback on the initial stage of this work. The author also acknowledges helpful discussion with Dr. Amrita Mukherjee from TIFR, Mumbai on the calculation of the persistent current.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request. *
Appendix A Results for higher generation system
We remark that, as our model demand a very complicated implementation scheme in an actual real-life experimental situation, it could be difficult to fabricate and do the experiment for a large system. However, to convince the readers, in this appendix, we show the results for a 5th generation system (Fig. 8) with $\mathcal{N}(5)=366$ sites as displayed in Fig. 9.
Figure 8: Schematic representation of the $5$th generation SPG fractal lattice geometry with hierarchical distribution of magnetic fluxes in the downward triangular plaquettes. Different colors represent the hierarchical flux distribution according to the order of hierarchy of the downward triangles (from smallest to the largest), viz., $\Phi_{1}$, $\Phi_{2}$, $\Phi_{3}$, $\Phi_{4}$, and $\Phi_{5}$.
Figure 9: Energy eigenvalue spectrum for a 5th generation SPG lattice structure (having $\mathcal{N}(5)=366$ sites) as a function of the hierarchical magnetic flux with values of the hierarchy parameter chosen to be same as in Fig. 2 and Fig. 3.
We found that, all our predictions from the previous generation systems (discussed in the main text) holds good for this relatively bigger system size also. It is worth mentioning that, as we have a deterministic disorder distribution of AB phases along different bonds for any given generation for our lattice model, there is no ready-made algorithm to generate the Hamiltonian for the system – one has do it by hand for each given generation.
Figure 10: Schematic representation of a SPG fractal structure with a uniform distribution of the magnetic flux in all the upward triangular loops (left panel) and its energy eigenvalue distribution (right panel).
In this study, we have considered the distribution of the magnetic flux in the downward triangular loops of the SPG fractal structure in such a way that it follows a special kind of hierarchical pattern. However, it is also interesting see what will happen if we put the magnetic flux in all the upward triangular loops of the SPG structure. This will impose a uniform Aharonov-Bohm phase factor along all the bonds in the SPG fractal structure (see Fig. 10). In this case, there will be no concept of hierarchical flux distribution as automatically imposed by the ramification pattern of the SPG fractal geometry. We have done the computation for this case also and the results for the energy eigenvalue spectrum for a finite generation SPG fractal structure is shown in Fig. 10.